# Arithmetic in Transformers Explained

### v9

@@ 4.2 Investigating Five-Digit Addition
L0H0 A2.ST A3.ST A1.ST A4.ST A4.SC A3.SC A2.SC A1.SC A0.SC
L0H1 A1.ST A0.ST A5.SA A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST
4.3 Investigating n-digit Addition models
To test whether this algorithm generalises beyond 5-digit addition, we conducted a comprehensive study of larger addition models. We trained 15 2-layer transformer models to handle addition problems ranging from 5- to 15-digits. These models used varying random seeds and optimizers, and included a sign token in their output format (e.g. 111111+222222=+0333333).
Figure 4: The 5- to 13-digit 2-layer 3-head addition models have very low loss and >99.999% accuracy. With more digits the model loss and accuracy is worse. Details in Tab.6.
Analysis Framework. To systematically analyse these models, we first created a declarative method for specifying each subtask within the addition process. Each subtask definition includes (1) the input digits it processes, (2) the answer digit it affects, (3) the complexity level it influences, and (4) expected PCA patterns and ablation test results. Using this declaration, we identified the nodes performing these subtasks across all models.
Since our framework lays out the addition algorithm in terms of its required subtasks., we can also verify whether a given model implements this algorithm by checking for the presence and proper sequencing of all required subtasks. For example, our addition algorithm specifies that the model must execute $A_{n}.ST$ subtasks for each question digit before the first answer token. Table 2 shows how an addition model fully implements all required subtasks in its nodes.
Using this framework, we found that all fifteen 2-layer addition models, regardless of their size, implemented the same core algorithm we identified in our 5-digit model. This consistent pattern suggests we have identified a robust and generalizable algorithm that naturally arises when transformer models are trained on addition in our setup where the digits are tokenized individually.
4.4 Training n-digit Mixed models
We trained 14 mixed models with various architectures, including 6- and 10-digit models with 2-3 layers and 3-4 attention heads. Each model was initialized using weights from a highly accurate addition-only model, specifically copying the weights into the first two layers and first three heads of the mixed model. During training, we used a curriculum of 80% subtraction and 20% addition examples, with deliberate enrichment of rare edge cases for both operations.
This initialization approach proved effective, with most models achieving high accuracy on both operations. We explored alternative training strategies, such as periodically ”freezing” the inserted attention heads or MLPs by restoring their original addition-model weights every 100 training steps. However, these more constrained approaches resulted in lower accuracy, suggesting that some flexibility in modifying the initialized weights was necessary for optimal performance.
The initialized mixed experimental models achieve >99.999% accuracy and contain the same subtasks, implying a common algorithm.
4.5 Investigating n-digit Mixed models
The mixed models revealed fascinating patterns in how they handled different types of arithmetic problems. Our hypothesis was that the model handled three distinct question classes: addition, positive-answer subtraction (e.g., 10009-10000=+000009), and negative-answer subtraction (e.g., 10009-20000=-009991).
Through ablation studies, we found that many nodes specialized in handling specific question classes, while others became ”polysemantic,” capable of processing multiple classes simultaneously (More details in Tab.3). Most notably, the majority of nodes initialized from the addition model became polysemantic during training. Rather than developing entirely new circuits for subtraction, the models adapted their addition circuits to handle both operations.
@@ 4.5 Investigating n-digit Mixed models
Positive-answer sub 70 73% 40 83%
Negative-answer sub 53 55% 29 60%
Defining Subtraction Subtasks. Subtraction requires two distinct computational paths based on the relative magnitude of the numbers. When D $\geq$ D’, as in 325 - 123 = +202, we perform direct subtraction with borrowing. When D $<$ D’, as in 325 - 456 = -131, we must instead compute the magnitude of the difference and mark it negative.
Hence, this computational difference necessitates separate frameworks for positive-answer (M) and negative-answer (N) subtraction, each with parallel operations to addition’s Base Add ($SA$) and Carry One ($SC$), which we define in Tab. 4. We also define complexity measures for subtraction (see App. G and  N).
Table 4: We define 2 “positive-answer subtraction” and 2 “negative-answer subtraction” subtasks that parallel the addition subtasks, Base Add SA and Carry One SC .
Name +ve Sub -ve Sub Definition
Base Diff MD ND $D_{n}$- $D^{\prime}_{n}$10
Borrow One MB NB $D_{n}$- $D^{\prime}_{n}$$<0$
Just as addition must handle ”cascading carry” cases (e.g., 99999+1=100000), subtraction must handle ”cascading borrow” cases. Consider 100000-1=099999, where a borrow cascades through multiple zeros. To handle such cases efficiently, we define a three-state classification $MT$ that parallels addition’s $ST$ :
$\begin{split}A_{n}.MT&=\underbrace{\text{TriCase}}_{(D_{n},D^{\prime}_{n})}=\\ &\begin{cases}MTN&\text{if }D_{n}<D^{\prime}_{n}\\ MT0&\text{if }D_{n}=D^{\prime}_{n}\\ MT1&\text{if }D_{n}>D^{\prime}_{n}\\ \end{cases}\end{split}$ (5)
For M questions, MTN is a definite borrow one, MT0 is a possible borrow one (depending on the results of other digit calculations) and MT1 is definitely not a borrow one. For N questions, the interpretation is the opposite.
Finally, paralleling the ”cascading carry one” $An.SV$ operation in addition, we define cascading borrow calculations $A_{n}.MV$ and $A_{n}.NV$ for positive and negative cases.
Evolution of Polysemantic Nodes. During training, the mixed model adapted the nodes it inherited from the addition model. Nodes that originally only performed Simple Addition ($SA$) developed the ability to simultaneously handle three distinct operations: addition ($SA$), positive-answer subtraction ($MD$), and negative-answer subtraction ($ND$). (Refer to Table 5.) We believe that this adaptation was likely because these three operations share a common structure: each maps a pair of digits (10 possibilities for $D_{n}$× 10 possibilities for $D^{\prime}_{n}$= 100 input cases) to a single-digit result (10 possible outputs). The model similarly adapted nodes that originally handled carry bits ($SC$) to also process borrow operations for both positive ($MB$) and negative ($NB$) subtraction cases.
Table 5: For mixed models, in later tokens, polysemantic attention heads simultaneously generate outputs for the three question classes addition S, M and N. Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token. The MLP layers then select the output appropriate for the class. In this sample map, from P16, the first 3 rows contain many polysemantic nodes, while the 4th row calculates the question class.
(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
L0H0 A4.MT A3.MT A4.ST A2.SC A1.SC A0.SC OPR
A4.GT A3.MT A4.MT A4.SC A3.SC A2.NB A1.MB A0.MB SGN
OPR A1.NB A0.NB
L0H1 A2.ST A1.ST A3.ST A0.ST A5.SA
A4.ST A2.MT A1.MT A3.GT A0.MT A5.MD
A2.GT A1.GT A0.GT A4.SA A3.SA A2.SA A1.SA A0.SA
L0H2 A5.ST A5.SA A4.MD A3.MD A2.MD A1.MD A0.MD
OPR A5.MD A4.ND A3.ND A2.ND A1.ND A0.ND
SGN A5.ND
L0H3 OPR OPR OPR OPR OPR OPR
SGN SGN SGN SGN SGN SGN
4.6 Mixed model question class detection
Mixed models must identify whether each input belongs to one of three classes: addition (S), positive-answer subtraction (M), or negative-answer subtraction (N). For addition questions (marked by ”+”), classification is immediate. For subtraction questions (marked by ”-”), the model must determine whether D $\geq$ D’ to distinguish between M and N classes. This classification must occur by the ”=” token since it determines whether the first answer token will be ”+” or ”-”.
The model can determine D $\geq$ D’ in two ways. First, it can use the $A_{n}.MT$ values already computed for subtraction. Alternatively, it can use a dedicated comparison circuit based on a new subtask:
$A_{n}.GT=\begin{cases}1&\text{if }D_{n}\geq D^{\prime}_{n},\\ 0&\text{otherwise}.\end{cases}$ (6)
For a 4-digit number, this comparison cascades through the digits: D $\geq$ D’ is true if either $A3.GT=1$, or if $A3.GT=0$ and $A2.GT=1$, and so on through $A0.GT$. We verify this circuit through ablation tests that show its disruption causes answer signs to flip.
We found all mixed models incorporate the $A_{n}.MT$ and $A_{n}.GT$ subtasks to determine whether the answer is postiive or negative. Typically (but not always), the same node computes both subtasks (see Table 2). Since both the $MT$ and GT approaches can validly distinguish between positive and negative outcomes, the model sometimes chooses one approach for some digits and the other approach for others.
A single node can implement multiple arithmetic operations (e.g., $A2.SA$, $A2.MD$, $A2.ND$) but has no intrinsic way of knowing whether the current question is addition, positive subtraction, or negative subtraction. Hence, we hypothesise that it outputs all three possible results into the residual stream. A different node must then select the correct result by identifying the question’s operation and required sign.
In practice, we observe that one or more attention heads attend to the question operator token (OPR) and the initial answer sign token (SGN), encoding the necessary information for the downstream MLP layers to pick which of the three outputs is correct. As Table 2 shows, each answer digit typically has one attention head that focuses on both the OPR and SGN tokens, thereby enabling the MLP to filter out the irrelevant outputs and retain the correct one.
@@ 5 Conclusion
Our analysis of 44 transformer models demonstrates their ability to learn highly accurate mathematical algorithms, achieving over 99.999% accuracy on addition tasks up to 13 digits. By investigating how models implement computations using attention heads and MLPs, we uncovered novel computational approaches distinct from traditional human calculation methods. Parameter transfer experiments revealed that features originally specialized for addition could be repurposed for subtraction, with many nodes becoming polysemantic.
@@
Impact Statement
Appendix A Appendix: Terminology
@@
  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+” or ”-” answer sign
  • S : Prefix for Addition. Think S for Sum. Aka ADD.
  • SA : Basic Add. An addition subtask. An.SA is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
  • SC : Carry One. An addition subtask. An.SC is defined as Dn + D’n >=>=> = 10. e.g. 5 + 7 gives True
  • SS : Make Sum 9. An addition subtask. An.SS is defined as Dn + D’n == 9. e.g. 5 + 7 gives False
  • ST : TriCase. An addition subtask. Refer paper 2 for details
  • ST8, ST9, ST10 : Outputs of the ST TriCase subtask.
@@
  • : Basic Difference. A subtraction subtask. An.MD is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • : Borrow One. A positive-answer subtraction subtask. An.MB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • MZ : Make Zero. A positive-answer subtraction subtask. An.MZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • MT : TriCase. A positive-answer subtraction subtask.
  • MT1, MT0, MT-1 : Outputs of the MT TriCase subtask.
@@
  • ND : Basic Difference. A negative-answer subtraction subtask. An.ND is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • NB : Borrow One. A negative-answer subtraction subtask. An.NB is defined as Dn - D’n $<$ 0. e.g. 5 - 7 gives True
  • NZ : Make Zero. A negative-answer subtraction subtask. An.NZ is defined as Dn - D’n == 0. e.g. 5 - 5 gives True
  • NT : TriCase. A negative-answer subtraction subtask.
  • GT : Greater Than. A (positive-answer or negative-answer) subtraction subtask. An.GT is defined as Dn $>$ D’n. e.g. 3 $>$ 5 gives False
@@
Appendix B Appendix: Model Configuration
@@
  • n_digits = 5, 6 or 10: Number of digits in the question.
Each digit is represented as a separate token. (Liu & Low, 2023) state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”. The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
Training uses a new batch of data each step (aka Infinite Training Data) to minimise memorisation. Depending on the configuration, each training run processes 1 to 4 million training datums. For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems.
Addition and subtraction include rare edge cases. For example, the SS cascades (e.g. 44445+55555=100000, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare. The data generator was enhanced to increase the frequency of all known edges cases. This lead to lower model loss.
Validation test data covering all edge cases was manually constructed. These test cases are not used during training.
The Colabs will be made available on publication.
Appendix C Appendix: Model Loss
The model defaults to batch size = 64, learning rate = 0.00008 and weight decay = 0.1. The loss function is simple:
In our experimental models, the number of digits in the question varies from 5 to 13, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4. Each experimental model’s loss is detailed in Tabs. 6 and  7.
Appendix D Appendix: Addition Model Shape
While we wanted a very low loss addition model, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model. Here are the things we tried to reduce loss that didn’t work:
  • Increasing the frequency of hard (cascading SS ) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
@@ Appendix D Appendix: Addition Model Shape
Appendix E Appendix: Experimental models
Forty-four models were trained and analyzed (Refer Tabs. 6 and  7). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke in the quanta_maths_* collections to support further research in AI Safety.
@@ Appendix E Appendix: Experimental models
Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Heads used MLPs used
5 1 3 30K 372001 9.4e-2 12621 15 6
5 2 3 15K 372001 1.6e-8 0 30 16
5 2 3 40K 372001 2.0e-9 0 22 15
6 2 3 15K 372001 1.7e-8 2 31 17
6 2 3 20K 173289 1.5e-8 0 28 17
6 2 3 20K 572091 7.0e-9 0 35 17
6 2 3 40K 372001 2.0e-9 0 29 17
7 2 3 45K 173289 3.0e-9 0 31 20
8 2 3 45K 173289 3.0e-9 0 35 22
9 2 3 45K 173289 3.0e-9 0 54 27
10 2 3 40K 572091 7.0e-9 0 44 28
11 2 3 50K 173289 8.0e-9 2 56 29
12 2 3 50K 173289 5.0e-9 3 50 33
13 2 3 50K 173289 6.3e-8 1 66 31
14 2 3 60K 173289 5.5e-6 199 68 35
15 2 3 80K 572091 8.6e-8 10 93 58
@@
Num Digits Num Layers Num Heads Train Steps Train Seed Train loss Addition Fails / M Subtract Fails / M Heads used MLPs used
@@
6 4 4 40K 372001 1.7e-8 3 8 51 30
Mixed models initialized with add model. Reset useful heads & MLPs every 100 steps
6 4 3 40K 372001 3.0e-4 17 3120 115 53
@@
  • A “features.json” file containing maths-specific ”feature” facts learnt about the model by the Colab e.g. P18L0H0 performs the A3.SC subtask.
Appendix F Appendix: TriAdd Implementation
TriAdd transfers data from $A_{n-1}$to $A_{n}$by integrating the values of $A_{n-1}.ST$ and $A_{n}.ST$ This function can be represented as nine bigram mappings with three possible outputs. (Refer Tab.8.)
Table 8: $A_{n}$.TriAdd can be calculated from $A_{n}.ST$ and $A_{n-1}.ST$ through nine bigram mappings and yielding the three distinct outputs ST8, ST9 and ST10
$A_{n}.ST$ $A_{n}.ST$ $A_{n}.ST$
$A_{n-1}.ST$ = ST8 = ST9 = ST10
ST8 ST8 ST9 ST10
ST9 ST8 ST9 ST10
 ST10 ST8 * ST10 ST10
Note that in the case $A_{n}.ST$ = ST9 and $A_{n-1}.ST$ = ST10, the answer is indeterminate. The result could be ST8 or ST9 but importantly it can not be ST10. We choose to use ST8 in our definition, but ST9 would work just as well.
Appendix G Appendix: Complexity

### v10

@@ 4.2.2 Addition model results
L0H0 ST2 ST3 ST1 ST4 SC4 SC3 SC2 SC1 SC0
L0H1 ST1 ST0 SA5 SA4 SA3 SA2 SA1 SA0
L0H2 ST5
4.2.3 Mixed model results
Testing our accurate ”mixed” (both addition and subtraction) models, we find all models contain the subtasks required by our addition algorithm, and satisfy the algorithmic requirements.
The mixed models have the same types of variability that the addition models have.
Evolution of Polysemantic Nodes. Zhang et al. (2024) identified symmetries between addition and subtraction circuits, with the same attention heads handling both operations. We expand on this finding.
Consider the mixed models that we initialized with accurate, smaller addition models before training. The majority of inherited nodes became polysemantic during training. Rather than developing entirely new circuits for subtraction, the models adapted their addition circuits to handle both operations. For example, a node that originally only performed Base Add ($SA$ ) often learns to handle three subtasks: addition ($SA$ ), positive-answer subtraction diff ($MD$ ), and negative-answer subtraction diff ($ND$ ). (Refer to Tab. 2.)
Table 2: For this mixed model, in the last 5 tokens, polysemantic attention heads simultaneously generate outputs for the three question classes (S, M and N). Other heads calculate the question class by attending to the question operation (OPR) token and the answer sign (SGN) token. We assume the MLP layers then select the output appropriate for the class.
(P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17) (P18) (P19) (P20)
D’3 D’2 D’1 D’0 = A7 A6 A5 A4 A3 A2 A1
L0H0 MT4 MT3 ST4 SC2 SC1 SC0 OPR
MT3 MT4 SC4 SC3 NB2 MB1 MB0 SGN
OPR NB1 NB0
L0H1 ST4 ST2 ST1 ST3 ST0 SA5
MT2 MT1 MT0 MD5 SA4 SA3 SA2 SA1 SA0
L0H2 ST5 SA4 MD4 MD3 MD2 MD1 MD0
OPR ND5 ND4 ND3 ND2 ND1 ND0
SGN
L0H3 OPR OPR OPR OPR OPR OPR
SGN SGN SGN SGN SGN SGN
We believe that this adaptation was likely because these three operations share a common structure: each maps a pair of digits (10 possibilities for $D_{n}$× 10 possibilities for $D^{\prime}_{n}$= 100 input cases) to a single-digit result (10 possible outputs). Similarly, some models adapted nodes that originally handled carry bits ($ST$) to also process borrow bit operations ($MT$).
@@ 4.2.3 Mixed model results
Positive-answer subtraction 70 73% 40 83%
Negative-answer subtraction 53 55% 29 60%
@@ 5 Conclusion
Our analysis of 49 transformer models demonstrate their ability to learn highly accurate mathematical algorithms, achieving over 99.999% accuracy on addition tasks up to 15 digits. By investigating how models implement computations using attention heads and MLPs, we uncovered novel computational approaches distinct from traditional human calculation methods. Parameter transfer experiments revealed that features originally specialized for addition could be repurposed for subtraction, with many nodes becoming polysemantic.
@@
6 Limitations and Future Work
While we identify and test the functional role of each component in the mixed model algorithm, we do not analyze the specific data representations and transformations used by polysemantic nodes, SGN nodes, and OPR nodes in the residual stream, nor the detailed mechanisms by which MLP layers process this information. Although our algorithmic-level analysis may be more generalizable across models, this approach leaves gaps in understanding how the identified components actually encode and manipulate information.
Our automated framework for discovering algorithm subtasks in models, while instrumental in accelerating our research, has limitations. Some aspects are specific to our math models. Some aspects are generic and have already been used in the published paper by  Harrasse et al. (2025).
Fully reverse engineering neural networks faces significant challenges, including distributed parallel computations difficult to trace to specific components. We made progress on an alternative approach: a declarative language describing algorithms in terms of subtasks and a framework to test these descriptions against models. However, more work is needed to refine this method and quantity our certainty when comparing evidence gathered from models against an algorithmic description.
Our models may be useful to researchers investigating the universality of representations / circuits across models.
Appendix A Appendix: Terminology and Abbreviations
@@
  • An : nth token in the answer. Zero-based. A0 is the units value. The highest token is the “+” or “-” answer sign
  • S : Prefix for Addition. Think S for Sum.
  • SA : Base Add. An addition subtask. $SA_{n}$ is defined as (Dn + D’n) % 10. e.g. 5 + 7 gives 2
  • SC : Carry One. An addition subtask. $SC_{n}$ is defined as Dn + D’n >=>= 10. e.g. 5 + 7 gives True
  • ST : TriCase. An addition subtask. Refer paper body for details
@@
  • MD: Basic Difference. A subtraction subtask. $MD_{n}$ is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • MB: Borrow One. A positive-answer subtraction subtask. $MB_{n}$ is defined as Dn - D’n << 0. e.g. 5 - 7 gives True
@@
  • ND : Basic Difference. A negative-answer subtraction subtask. $ND_{n}$ is defined as (Dn - D’n) % 10. e.g. 3 - 7 gives 6
  • NB : Borrow One. A negative-answer subtraction subtask. $NB_{n}$ is defined as Dn - D’n << 0. e.g. 5 - 7 gives True
@@
Appendix B Appendix: Ethics Statement
Appendix C Appendix: Vocabulary
Each digit is represented as a separate token. [Liu and Low, 2023] state that LLaMa’s “remarkable arithmetic ability … is mainly atributed to LLaMA’s consistent tokenization of numbers”. The model’s vocabulary contains 14 tokens ( 0, .., 9, +, -, =, *, / ) to enable this and planned future investigations.
Appendix D Appendix: Training Data
Training uses a new batch of data each step (aka Infinite Training Data) to minimize memorization. Depending on the configuration, each training run processes 1 to 4 million training datums. For the 5-digit addition problem there are 100,000 squared (that is 10 billion) possible questions. So the training data is much less than 1% of the possible problems.
Addition and subtraction include rare edge cases. For example, these cascades (e.g. 55555+44446=100001, 54321+45679=1000000, 44450+55550=10000, 1234+8769=10003) are exceedingly rare. The data generator was enhanced to increase the frequency of all known edges cases. This leads to lower model loss.
We enriched 60% of training data based on these edge cases (leaving the other 40% of training data random): (1) The initial model failed at cascading carry ones in addition. To rectify this, we randomly selected one operand and modified a random subset of digits in that operand to make the selected digit-position sum to 9, increasing the likelihood of a cascading carry one. (2) The initial model was worse at the subtraction task when the answer was negative, so we added 1 to each second operand digit (that was 8 or less) to increase the frequency of negative answers. (3) When the operands were identical, the initial model predicted -000000 instead of +000000. We increased the frequency of this case. For example, for 6 digit questions, we increased the frequency from 0.0001% to 0.6%.
Appendix E Appendix: Surveying LLM Addition Capability
LLM Gateways provide access to hundreds of LLMs. We used the Martian Gateway ( https://app.withmartian.com/ ) to quickly and cheaply access 193 LLMs ( https://withmartian.github.io/llm-adapters/ ) ranging in size from 1 to 405 billion parameters. We tested their ability to perform addition accurately on the test questions containing cascading carry ones shown in Tab. 4. A model’s score represents the longest question the model succeeded on before it failed a question. The results are shown in Fig. 2. Fig. 8 shows the size of the models seems uncorrelated to its ability to perform these questions.
Table 4: Addition prompts used to test LLMs’ ability to handle the cascading carry one use case.
Test Prompt Correct Answer
Answer concisely: 6+5= 11
Answer concisely: 19+87= 106
Answer concisely: 774+229= 1003
Answer concisely: 6587+3416= 10003
Answer concisely: 22605+77398= 100003
Answer concisely: 532847+467159= 1000006
Answer concisely: 5613709+4386294= 10000003
Answer concisely: 72582383+27417619= 100000002
Answer concisely: 206727644+793272359= 1000000003
Answer concisely: 7580116456+2419883549= 10000000005
Answer concisely: 52449010267+47550989737= 100000000004
Answer concisely: 888522030597+111477969406= 1000000000003
Figure 8: This scatter plot shows each model’s addition and subtraction score against the model output token cost (an approximate metric for the model size). The maximum possible score is 15.
At time of writing, 4.76% of models scored 10 or more (specifically nousresearch/hermes-3-llama-3.1-405b, amazon/nova-pro-v1, anthropic/claude-opus-4-1-20250805, mistralai/magistral-medium-2506, openai/gpt-4o-mini-search-preview, openai/gpt-4o-search-preview, openai/o4-mini, qwen/qwen3-coder and tencent/hunyuan-a13b-instruct). Some of these models explicitly acknowledge the use of external tools. For others it is not obvious whether they use tools. The three Chat GPT 5 models tested scored 4, 4 and 7. After publication, we will make public the Jupyter notebook used to perform the test. It takes less than 20 minutes and costs less than 1 USD to run.
Appendix F Appendix: Training Setup
@@
  • n_digits = 5 .. 15: Number of digits in the question.
The Colabs are available at https://github.com/PhilipQuirke/quanta_maths
While we wanted a very low loss models, we also wanted to keep the model compact - intuiting that a smaller model would be easier to understand than a large model. Here are the things we tried to reduce loss that didn’t work:
  • Increasing the frequency of hard (cascading carry one) examples in the training data so the model has more hard examples to learn from. This improved training speed but did not reduce loss.
@@
  • Inserting “+” in the answer format (e.g. “12345+22222=034567” becomes “12345+22222=+034567” had no impact on accuracy or the algorithm.
@@
Appendix G Appendix: Model Loss
The model defaults to batch size = 64. The loss function is simple:
In our experimental models, the number of digits in the question varies from 5 to 15, the number of layers varies from 1 to 4, the number of heads varies from 3 to 4. Each experimental model’s loss is detailed in Tabs. 5 and  6.
During training, the models use an AdamW optimizer with a learning rate (LR) of 0.00008, weight decay of 0.1 and betas of (0.9, 0.98). During the first 5th of training the LR is linearly increased from 0.01 * LR to LR to help stabilize training. During the remainder of training, cosine annealing is used.
Appendix H Appendix: Experimental models
Forty-nine models were trained and analyzed (Refer Tabs. 5 and  6). The models and output of analysis are available on HuggingFace at https://huggingface.co/PhilipQuirke in the quanta_maths_* collections to support further research in AI Safety.
@@ Appendix H Appendix: Experimental models
Digits Layers Heads Train Steps Train Seed Train loss Add Fails / M Heads used MLPs used Clopper-Pearson 95% Interval
5 1 3 30K 372001 9.4e-2 12621 15 6 [1.24e-2, 1.28e-2]
5 2 3 15K 372001 1.6e-8 0 30 16 [0.00e+0, 3.69e-6]
5 2 3 40K 372001 2.0e-9 0 22 15 [0.00e+0, 3.69e-6]
6 2 3 15K 372001 1.7e-8 2 31 17 [2.42e-7, 7.22e-6]
6 2 3 20K 173289 1.5e-8 0 28 17 [0.00e+0, 3.69e-6]
6 2 3 20K 572091 7.0e-9 0 35 17 [0.00e+0, 3.69e-6]
6 2 3 40K 372001 2.0e-9 0 29 17 [0.00e+0, 3.69e-6]
7 2 3 45K 173289 3.0e-9 0 31 20 [0.00e+0, 3.69e-6]
8 2 3 45K 173289 3.0e-9 0 35 22 [0.00e+0, 3.69e-6]
9 2 3 45K 173289 3.0e-9 0 54 27 [0.00e+0, 3.69e-6]
10 2 3 40K 572091 7.0e-9 0 44 28 [0.00e+0, 3.69e-6]
11 2 3 50K 173289 8.0e-9 2 56 29 [2.42e-7, 7.22e-6]
12 2 3 50K 173289 5.0e-9 3 50 33 [4.25e-7, 7.88e-6]
13 2 3 50K 173289 6.3e-8 1 66 31 [2.53e-8, 5.57e-6]
14 2 3 60K 173289 5.5e-6 199 68 35 [1.74e-4, 2.00e-4]
15 2 3 80K 572091 8.6e-8 10 93 58 [4.71e-6, 1.67e-5]
@@
Digits Layers Heads Train Steps Train Seed Train loss Add Fails / M Sub Fails / M Heads used MLPs used Clopper-Pearson 95% Interval
@@
6 4 4 40K 372001 1.7e-8 3 8 51 30 [5.64e-6, 2.15e-5]
Mixed models inited with add model. Reset useful hds & MLPs every 100 steps
6 4 3 40K 372001 3.0e-4 17 3120 115 53 [3.06e-3, 3.30e-3]
@@
  • A “features.json” file containing maths-specific “feature” facts learnt about the model by the Colab e.g. P18L0H0 performs the SC3 subtask.
Table 7: For a sample model, all nodes used in predictions are shown by token position (horizontally) and model layer (vertically), detailing the answer digits they impact. Here, the attention heads in token position P10 labelled A5..3 help predict the answer digits A3, A4 and A5. For all addition and mixed models studied, before the “=” token, each node often calculates data used to predict multiple answer digits. After the “=” token, all nodes in a given token position are used to predict a single answer digit.
(P6) (P9) (P10) (P11) (P12) (P13) (P14) (P15) (P16) (P17)
D’4 D’1 D’0 = + A5 A4 A3 A2 A1
L0H0 A4 A2 A1
L0H1 A5 A5..3 A5 A3 A0
L0H2 A5 A5..1
L0MLP A5..2 A4 A3 A2 A1 A0
L1H0
L1H2 A4 A3 A2
L1MLP A1 A0
Appendix I Appendix: Complexity